OpenStudy (anonymous):

A student solved 3x + 4y = 12 for y. Her work is at the left. What error did the student make?

OpenStudy (anonymous):

3x + 4y = 12 4y = -3x+12 y = -3/4 + 3

OpenStudy (anonymous):

that is the correct answer

OpenStudy (anonymous):

dont forget the x after -3/4 so it would be y = -3/4x + 3

OpenStudy (anonymous):

thanks